The equation of the plane passing through the point (1, –2, 4) and perpendicular to the line, whose direction ratios are 2, 4, –1 is


1
Expert's answer
2021-12-16T07:30:23-0500

Equation of a plane perpendicular to a given vector "\vec n=\langle2, 4, -1\rangle" and passing through a given point "(1, \u20132, 4)" is


"2(x-1)+4(y-(-2))+(-1)(z-4)=0"

"2x-2+4y+8-z+4=0"

The equation of the plane passing through the point (1, –2, 4) and perpendicular to the line, whose direction ratios are 2, 4, –1 is

"2x+4y-z+10=0"